# Gen-C: Populating Virtual Worlds with Generative Crowds

### v1

@@ Abstract
Over the past two decades, researchers have made significant advancements in simulating human crowds, yet these efforts largely focus on low-level tasks like collision avoidance and a narrow range of behaviors such as path following and flocking. However, creating compelling crowd scenes demands more than just functional movement—it requires capturing high-level interactions between agents, their environment, and each other over time. To address this issue, we introduce Gen-C, a generative model to automate the task of authoring high-level crowd behaviors. Gen-C bypasses the labor-intensive and challenging task of collecting and annotating real crowd video data by leveraging a large language model (LLM) to generate a limited set of crowd scenarios, which are subsequently expanded and generalized through simulations to construct time-expanded graphs that model the actions and interactions of virtual agents. Our method employs two Variational Graph Auto-Encoders guided by a condition prior network: one dedicated to learning a latent space for graph structures (agent interactions) and the other for node features (agent actions and navigation). This setup enables the flexible generation of dynamic crowd interactions. The trained model can be conditioned on natural language, empowering users to synthesize novel crowd behaviors from text descriptions. We demonstrate the effectiveness of our approach in two scenarios, a University Campus and a Train Station, showcasing its potential for populating diverse virtual environments with agents exhibiting varied and dynamic behaviors that reflect complex interactions and high-level decision-making patterns.
@@ 1 Introduction
To facilitate digital human creation, the field of computer graphics has experienced a dramatic increase in the number of tools, approaches and algorithms focusing on creating compelling crowd motions [1, 2]. Currently, several game engines and companies specializing in offline modeling and rendering offer tools to author crowd simulations [3, 4, 5, 6]. Nevertheless, a lot of manual work is still needed by the animators and level designers along with deep familiarity with the underlying tools in the production workflow in order to obtain desired simulation results. While a number of automated solutions to crowd simulation have been proposed over the past decade, they mostly focus on low-level tasks like resolving collisions between agents and a narrow range of behaviors such as steering and rouping [7]. In contrast, real humans are capable of exhibiting high-level behaviors of enormous complexity and diversity based on their own goals and desires and the affordances from the environment. For example, we stop on the street to briefly talk to friends, browse through shopping windows, buy a ticket before heading to the train platform, etc. Currently, such behaviors are either absent from existing simulators resulting in "boring" and repetitive simulations or are scripted which makes them difficult to scale.
@@
To break free from the dependency on real crowd data, we propose to automatically synthesize interaction datasets by leveraging the power of large language models (LLMs) such as the GPT framework [12]. As querying LLM APIs can become prohibitively expensive, we use them to create a few indicative crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently expand and generalize through simulations to construct corresponding crowd scenario graphs. Given such graphs, a generative model is trained to learn a latent space of agent behaviors and interactions and automate the synthesis of new graphs. As most existing graph embedding algorithms are mainly developed for static graphs, they cannot capture the complex and dynamic nature of a crowd scenario graph. To that end, we propose a novel model architecture based on variational graph autoencoders (VGAEs) [13] that can learn how agents interact with each other and how they behave with respect to their environment. The trained model can be conditioned on natural language, allowing users to automatically author scenes with diversified crowds from text input.
Our approach is inspired by recent trends in generative AI. We note that a lot of work in computer graphics and vision has applied generative models for text-driven human motion synthesis [14] with notable success. Despite being capable of synthesizing diverse behaviors, such works typically focus on individual humanoids.
More closely to our domain, generative models have also been explored to steer groups of agents [15] as well as to synthesize crowd trajectories [16] by leveraging guided diffusion. Our work is complementary to such approaches, as we seek to learn the high-level behaviors of crowds that can facilitate the generation of such trajectories. By leveraging a generative model, we strive for an approach that can generalize to different settings and number of agents, allowing characters to exhibit an array of high-level behaviors.
Overall, we propose:
@@
    Gen-C, a framework for text-guided authoring of high-level crowd behaviors that
@@
    breaks free from the dependency on real-world crowd data via the use of LLMs and simulations through
@@
    a text-conditioned graph generative model that relies on a dual VGAE architecture to encode the spatial and temporal evolution of agent behaviors, interactions, and movement.
@@ 2.1 Authoring Virtual Crowds
Controlling crowd simulations at a higher level is crucial to several scenarios since it allows users to easily and efficiently author agents’ behaviors according to their wishes. This requires intuitive tools which are highly dependent on what simulation aspect the users aim to control [2]. Early approaches to authoring large crowds included the use of pre-computed collision free trajectories in the form of crowd patches [17, 18] or crowd animations that could be connected together to form larger simulations  [19]. Others, devised user interactive tools to allow for the definition of navigation fields in pre-defined environments [20]. For authoring navigation behavior in crowd scenes, past literature heavily uses sketch-based interfaces [21, 22]. In contrast, authoring the animation and visualization aspects usually entails asset and template manipulation [23, 24]. Editing has also been widely explored in literature, with the more user-friendly systems incorporating manipulation handles. For example, Kwon et al. [25], propose deformation gestures as a post-processing tool for group motion editing. One of the most challenging aspects in authoring is controlling high-level behaviors such as describing agendas, desires and even personalities; these behaviors benefit the expressiveness and realism of simulations [26]. Complimentary to these approaches, we propose a text-driven generative approach to automatically author high-level crowd behaviors in specific environments, capitalizing on the implicit knowledge of LLMs for human decision making in different urban spaces.
@@ 2.2 Heterogeneity in Virtual Crowds
Generating behavioral diversity in virtual crowds has been addressed in the literature with a variety of approaches; early works use predefined rules for different behaviors to achieve such heterogeneity [27, 28]. Ren et al.  [29] incorporated diverse properties in groups of agents controllable via user constraints, while more generically several researchers [30, 31, 32] focused on behavior diversity in terms of personality traits; these expose some parameters to users to allow for control of agent behaviors. Data-driven approaches have also been prominent in the literature, where the crowd simulation model is implicitly defined by example data. Interestingly, early data-driven approaches were graph-based [33, 25], similar to early methods in the character animation literature [34]. These methods however remain limited to group navigation (e.g., flocking) and fail to reflect the variations of behaviors found in human crowds. The need to simulate behavior diversity found in real-life crowds, led to more sophisticated and practical approaches, including approaches that learn crowd models from trajectories extracted from real-world crowd data [35, 36, 37], including works that define secondary high-level actions for increased realism [38, 39], blending different crowd styles represented by the input data [21], and more recent solutions like deep reinforcement learning [40], and even in combination with imitation learning [16, 11]. Despite the significant improvements achieved by these methods, the results highly depend on the variability and amount of input data. Contrary to all of these works, Gen-C defines heterogeneity through the use of crowd scenario graphs that are automatically generated through LLM queries and a basic simulator allowing us to move away from the dependency on collecting real crowd data.
@@ 2.3 Large Language Models and Virtual Agents
Large language models have shown to be effective to domains beyond natural language processing such as task planning, embodied reasoning, autonomous agents, and floorplan design [41, 42, 43, 44]. Here, we explore an architecture that allows a LLM to generate a sequence of crowd behaviors over time and space that can then be generalized through simulations to create synthetic crowd data. We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors. Along these lines a lot of work has explored generative models for text-driven human motion generation tasks [14]. Despite impressive results, including generating scene-and-language conditioned motions [45, 46, 47, 48, 49], such solutions typically focus on generating animations for an individual character. But a human crowd is much more than that, consisting of multiple agents performing a sequence of behaviors while interacting with each other and the environment. Along these lines, the recent work of [15] uses a text-based diffusion model to steer crowds of agents but does not extend to other types of behaviors. Our work seek to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
@@ 2.4 Graph Generation
Besides our domain, learning graph representations is a fundamental task for many real-world applications [50]. While a lot of prior work has focused on learning static graphs with deep learning models such as VGAEs [13], recent techniques have explored learning of dynamic graphs where the graph structure itself changes as time progresses, including neural temporal point processes and attention mechanism for modeling time-dependent events [51], and deep architectures such as Temporal Graph Networks [52] and Spatio-Temporal Graph Convolutional Networks [53] that can capture temporal dependencies and predict future states of evolving graphs. While highly relevant to our work, we focus on crowd scenario graphs: highly dynamic graphs with complex features that capture the evolution of human decision-making. These graphs are more intricate than the time-evolving graphs studied in prior work, which are typically designed for tasks like link prediction, node classification, recommendation systems, and social network analysis. Closely related, conditional generation of graphs has also gained significant attention, with majority of works focusing on generating the graph structure only while the node and edge features are predetermined as compared to our approach. For example, NGG [54] uses conditioned latent diffusion, GraphVAE [55] integrates conditional codes for guided generation, and MOOD [56] guides a diffusion sampling process with property gradients. Sequential models like in [57] utilize conditions to initialize node and edge creation, while others incorporate conditional codes into message-passing schemes [58]. Complementary to these works , our Gen-C framework embeds both graph structures and node features into two distinct, prior-conditioned latent distributions, enabling their generation based solely on conditional text input.
@@ 3 Gen-C Overview
Preliminaries. We define a High-Level Scenario as the description of actions, interactions, and environment structure. Action is defined as a label from $\mathcal{A}ct=~{}${stand still, sit, wait, wander, queue, object interact, talk, meet, enter/exit, leave group, talk to phone, wave at, read, look at, carry}, that guides the agent’s behavior. When multiple agents perform a shared action, e.g. talk, we refer to that as an Interaction. Environment Structure defines the position, scale, and orientation of various locations; locations can be abstract, however we categorize them in: $\mathcal{L}oc=~{}${building, room, entrance, exhibit, furniture, outdoor area, item, service area}. To simplify our problem, we assume a predefined list of actions and location categories, and an action is always paired with a location, e.g. {talk, room}.
Gen-C Framework. Our proposed framework for populating virtual environments with generative crowds consists of three main components (Fig. 1): Synthetic Data Generation, Learning, and Crowd Scenarios Generation. To generate the synthetic data, we first use a LLM to create diverse crowd scenarios through targeted queries that capture context, environment placement, action probabilities, and movement probabilities. These scenarios are extended using a preliminary simulator to introduce randomness and generalize the data, which is then transformed into graph-like structures representing action sequences, relationships, and interactions. Given such graphs, we train two VGAE models: one learns to reconstruct the graph structures capturing agent interactions, while the other focuses on node features representing agent actions and navigation. A text-conditional network acts as a prior, integrating textual information to ensure the generated graphs align with given descriptions. The trained models can then synthesize novel crowd scenarios by conditionally sampling from the learned latent distributions, generating contextually aligned scenarios.
@@ 4.1 Querying Crowd Scenarios
To generate simulations that faithfully reflect crowd dynamics, we use a series of tailored LLM queries, each designed for a specific task. The process begins with a single sentence input $S_{in}$ describing a crowd scenario, which acts as the initial seed for the generation of a scenario (see top-left in Fig. 1 for an example input). Specifically, we select OpenAI’s gpt-4o [59] model and run four sequential queries as presented below:
Expanding the Scenario ($Q1$).
The first query requests the LLM to expand $S_{in}$ into a detailed paragraph $Q1_{par}$ that describes the scenario. This enriched description serves as the foundation for further steps as it includes individual agent behaviors and interactions, while also provides information about the environment context.
Defining the Environment ($Q2$).
Using the detailed description $Q1_{par}$ and a randomly selected environment size, the second query generates a meaningful environment layout $Q2_{env}$ specifying locations and areas relevant to the scenario. For example, the output might define places such as a “coffee shop”, “park”, or “entrance area”, along with their position, scale and orientation.
Assigning Location-Specific Actions ($Q3$).
For each generated location in $Q2_{env}$, the third query identifies a list of plausible actions from $\mathcal{A}ct$ that agents might perform in that area. For instance, in a “coffee shop,” actions like “queue,” “meet,” or “sit” are expected. The LLM takes as input the $Q1_{par}$ and $Q2_{env}$ and is requested to select the top-5 actions that are most suitable for each area and assign selection probabilities to them; we ensure the total probability sums to 1. We note that we allow multiple possible actions per location to promote diversity in the generated simulations.
Building Movement Plans ($Q4$).
The fourth query focuses on defining how agents move between different locations within the environment. This step is similar to $Q3$, however, assigns movement probabilities from one location to every other in the environment, creating a dynamic movement plan for the agents.
Finally, we combine the responses from all four queries and form a high-level crowd scenario $S$, generated by a high-level textual description $S_{in}$.
4.2 Collecting Individual Behaviors
After gathering multiple crowd scenarios from the LLM, we need a mechanism to utilize them so we can generate individual actions and interactions for each agent. Thus, we construct a simple simulator in Unity, where we parse each scenario $S$ and run multiple simulations.
For each LLM scenario, we build the environment as described by $Q2_{env}$. Then, we randomly spawn agents in the environment, either as individuals or in small groups of two or three (all agents in a group perform the same action). Initially, each agent or group, queries its current location action list and randomly selects a future action based on the assigned probabilities; each action has its own duration based on the context. After completing the action, the agent or group of agents selects the next location based on the movement probabilities of its current location. This process repeats in cycles, with data collected over a period of 2 to 3 minutes.
To introduce variability and prevent agents from consistently selecting actions and movements with the highest probabilities, we use a temperature parameter $\text{Temp}\in[.7,1]$ to amplify lower probabilities. Temp is randomly initialized and gradually increased to 1 by the end of the simulation.
Finally, for each agent, we record its action sequence throughout the simulation. Each entry in the sequence captures the selected action and its location. In case the current action is shared with others (interaction), we include the ID of other agent/s in the entry. This detailed record $Rec_{\alpha_{i}}$ for each agent ${\alpha_{i}}$, provides a comprehensive view of the agent’s behavior and interactions within the simulated environment over its entire lifespan.

### v2

@@ Abstract
Over the past two decades researchers have made significant advancements in simulating human crowds, yet these efforts largely focus on low-level tasks like collision avoidance and basic behaviors such as path following and flocking. Creating realistic crowd scenes, however, requires modeling richer, high-level interactions among agents, their environment, and each other over time. We introduce Gen-C, a graph-based generative model that shapes high-level agent actions and interactions. To bypass the labor-intensive and challenging task of collecting and annotating real crowd video data, we leverage a Large Language Model (LLM) to generate a dataset of crowd scenarios. Each scenario is then represented as a set of time-expanded graphs modeling agent actions and interactions. Gen-C employs a dual Variational Graph Autoencoder (VGAE) architecture to jointly learn graph structure (agent interactions) and features (agent actions), enabling the population of virtual scenes conditioned on natural language. We demonstrate the effectiveness of our approach across diverse scenarios, including a University Campus and a Train Station, showcasing its ability to populate virtual environments with agents that exhibit dynamic behaviors, complex interactions, and high-level decision-making.
@@ 1 Introduction
To facilitate digital human creation, the field of computer graphics has experienced a dramatic increase in the number of tools, approaches and algorithms focusing on creating compelling crowd motions [1, 2]. Currently, several game engines and companies specializing in offline modeling and rendering offer tools to author crowd simulations [3, 4, 5, 6]. Nevertheless, a lot of manual work is still needed by the animators and level designers along with deep familiarity with the underlying tools in the production workflow in order to obtain desired simulation results. While a number of automated solutions to crowd simulation have been proposed over the past decade, they mostly focus on low-level tasks like resolving collisions between agents and a narrow range of behaviors such as steering and grouping [7]. In contrast, real humans are capable of exhibiting high-level behaviors of enormous complexity and diversity based on their own goals and desires and the affordances from the environment. For example, we stop on the street to briefly talk to friends, browse through shopping windows, buy a ticket before heading to the train platform, etc. Currently, such behaviors are either absent from existing simulators resulting in "boring" and repetitive simulations or are manually scripted which makes them difficult to scale.
@@
To break free from the dependency on real crowd data, we propose to automatically synthesize interaction datasets by leveraging the power of Large Language Models (LLMs) such as the GPT framework [12]. As querying LLM APIs can become prohibitively expensive, we use them to create a few indicative synthetic crowd scenarios that describe agent-agent and agent-environment interactions for given input scenes, which we subsequently represent as crowd scenario graphs. Given such graphs, a generative model is trained to learn a latent space of agent behaviors and interactions, that automates the synthesis of new graphs (crowd scenarios). As most existing graph embedding algorithms are mainly developed for static graphs, they cannot capture the complex and dynamic nature of a crowd scenario graph. To that end, we propose a novel model architecture based on dual Variational Graph Autoencoders (VGAEs) [13] that can learn how agents interact with each other and how they behave with respect to their environment. The trained model can be conditioned mainly on natural language, allowing users to automatically author scenes with diversified crowds from text input.
Our approach is inspired by recent trends in generative AI. We note that a lot of work in computer graphics and vision has applied generative models for text-driven human motion synthesis [14] with notable success. Despite being capable of synthesizing diverse behaviors, such works typically focus on individual humanoids. More closely to our domain, generative models have also been explored to steer groups of agents [15] as well as to synthesize crowd trajectories [16] by leveraging guided diffusion. Our work is complementary to such approaches, as we seek to learn the high-level behaviors of crowds that can facilitate the generation of such trajectories. By leveraging a generative model, we strive for an approach that can generalize to different settings and number of agents, allowing characters to exhibit an array of high-level behaviors. Overall, we propose:
@@
    Gen-C, a framework for text-guided authoring of high-level crowd behaviors, that
@@
    breaks free from the dependency on real-world crowd data via the use of LLMs and simulations, through
@@
    a text-conditioned graph generative model which relies on a dual-VGAE architecture to encode the spatial and temporal evolution of agent behaviors, interactions, and movement.
@@ 2.1 Authoring Virtual Crowds
Controlling crowd simulations at a higher level is crucial to several scenarios since it allows users to easily and efficiently author agents’ behaviors according to their wishes. This requires intuitive tools which are highly dependent on what simulation aspect the users aim to control [2]. Early approaches to authoring large crowds included the use of pre-computed collision free trajectories in the form of crowd patches [17, 18] or crowd animations that could be connected together to form larger simulations  [19]. Others devised user interactive tools to allow for the definition of navigation fields in pre-defined environments [20]. For authoring navigation behavior in crowd scenes, past literature heavily uses sketch-based interfaces [21, 22]. In contrast, authoring the animation and visualization aspects usually entails asset and template manipulation [23, 24]. Editing has also been widely explored in literature, with the more user-friendly systems incorporating manipulation handles. For example, Kwon et al. [25] propose deformation gestures as a post-processing tool for group motion editing. One of the most challenging aspects in authoring is controlling high-level behaviors such as describing agendas, desires and even personalities; these behaviors benefit the expressiveness and realism of simulations [26]. Complimentary to these approaches, we propose a text-driven generative approach to automatically author high-level crowd behaviors in specific environments, capitalizing on the implicit knowledge of LLMs for human decision making in different urban spaces.
@@ 2.2 Heterogeneity in Virtual Crowds
Generating behavioral diversity in virtual crowds has been addressed in the literature with a variety of approaches; early works use predefined rules for different behaviors to achieve such heterogeneity [27, 28]. Ren et al. [29] incorporated diverse properties in groups of agents controllable via user constraints, while more generically several researchers [30, 31, 32] focused on behavior diversity in terms of personality traits; these expose some parameters to users to allow for control of agent behaviors. Data-driven approaches have also been prominent in the literature, where the crowd simulation model is implicitly defined by example data. Interestingly, early data-driven approaches were graph-based [33, 25, 34], similar to early methods in the character animation literature [35]. These methods however remain limited to group navigation (e.g., flocking) and fail to reflect the variations of behaviors found in human crowds. The need to simulate the behavioral diversity found in real-life crowds has led to more sophisticated and practical approaches. These include methods that learn crowd models from trajectories extracted from real-world crowd data [36, 37, 34], as well as works that define secondary high-level actions to increase realism [38, 39]. Additional approaches involve blending different crowd styles represented in the input data [21], and more recent solutions make use of deep reinforcement learning [40, 41]. Some even combine reinforcement learning with imitation learning [16, 11].
Despite the significant improvements achieved by these methods, the results highly depend on the variability and amount of input data. Contrary to all of these works, Gen-C defines heterogeneity through the use of crowd scenario graphs that are automatically generated through LLM queries allowing us to move away from the dependency on collecting real crowd data.
@@ 2.3 Large Language Models and Virtual Agents
Large language models have shown to be effective to domains beyond natural language processing such as task planning, embodied reasoning, autonomous agents, and floor plan design [42, 43, 44, 45]. Here, we explore an architecture that allows an LLM to generate a sequence of crowd behaviors, over time and space, that forms synthetic crowd datasets. We further exploit the power of generative NLP models to enable text-guided synthesis of high-level crowd behaviors. Along these lines a lot of work has explored generative models for text-driven human motion generation tasks [14]. Despite impressive results, including generating scene-and-language conditioned motions [46, 47, 48, 49, 50], such solutions typically focus on generating animations for an individual character. But a human crowd is much more than that, consisting of multiple agents performing a sequence of behaviors while interacting with each other and the environment. Along these lines, the recent work of [15] uses a text-based diffusion model to steer crowds of agents but does not extend to other types of behaviors. Our work seeks to learn high-level behaviors captured via crowd scenario graphs that can complement the low-level steering capabilities of such agents.
@@ 2.4 Graph Generation
Besides our domain, learning graph representations is a fundamental task for many real-world applications [51]. While a lot of prior work has focused on learning static graphs with deep learning models such as VGAEs [13], recent techniques have explored learning of dynamic graphs where the graph structure itself changes as time progresses, including neural temporal point processes and attention mechanism for modeling time-dependent events [52], and deep architectures such as Temporal Graph Networks [53] and Spatio-Temporal Graph Convolutional Networks [54] that can capture temporal dependencies and predict future states of evolving graphs. While highly relevant to our work, we focus on crowd scenario graphs: highly dynamic graphs with complex features that capture the evolution of human decision-making. These graphs are more intricate than the time-evolving graphs studied in prior work, which are typically designed for tasks like link prediction, node classification, recommendation systems, and social network analysis. Closely related, conditional generation of graphs has also gained significant attention, with majority of works focusing on generating the graph structure only while the node and edge features are predetermined as compared to our approach. For example, NGG [55] uses conditioned latent diffusion, GraphVAE [56] integrates conditional codes for guided generation, and MOOD [57] guides a diffusion sampling process with property gradients. Sequential models like in [58] utilize conditions to initialize node and edge creation, while others incorporate conditional codes into message-passing schemes [59]. Complementary to these works, Gen-C framework embeds both graph structures and node features into two distinct, prior-conditioned latent distributions, enabling their generation based solely on conditional text input.
@@ 3 Gen-C Overview
Preliminaries. We define a High-Level Scenario as the description of actions, interactions, and environment structure. Action is defined as a label from $\mathcal{A}ct=$~{wait, sit, wander, queue, enter/exit, wave at, discuss, meet, object interact, talk to phone, read, look at, carry}, that guides the agent’s behavior. When multiple agents perform a shared action, e.g. talk, we refer to that as an Interaction. Environment Structure defines the position, scale, and orientation of various locations; locations can be abstract, however we categorize them in: $\mathcal{L}oc=$~{building, room, entrance, exhibit, furniture, outdoor area, item, service area}. To simplify our problem, we assume a predefined list of actions and location categories, and an action is always paired with a location, e.g. {talk, room}.
Gen-C Framework. Our proposed framework for populating virtual environments with generative crowds consists of three main components (Fig. 1): Synthetic Data Generation, Learning, and Crowd Scenarios Generation. To generate the synthetic data, we first use a LLM to create diverse crowd scenarios through targeted queries that capture the actions and interactions of individuals and the environment placement. These scenarios are transformed into graph-like structures representing action sequences, relationships, and interactions. Given such graphs, we train two VGAE models: one learns to reconstruct the graph structures capturing agent interactions, while the other focuses on node features representing agent actions and navigation. A text-conditional network acts as a prior, integrating textual information to ensure the generated graphs align with the given descriptions. The trained models can then synthesize novel crowd scenarios by conditionally sampling from the learned latent distributions, generating contextually aligned scenarios.
@@ 4.1 Querying Crowd Scenarios
To generate simulations that faithfully reflect crowd dynamics, we use two tailored LLM queries, each designed for a specific task. The process begins with a single sentence input $S_{in}$ describing a crowd scenario, which acts as the initial seed for the generation of a scenario (see top-left in Fig. 1 for an example input). Specifically, we select OpenAI’s gpt-4.1 [60] model and run two sequential queries as presented below:
Generating the Environment ($Q1$). The first query requests the LLM to generate a meaningful environment layout, using $S_{in}$ as input, specifying locations and areas relevant to the scenario. For example, the output might define places such as a “coffee shop”, “park”, or “entrance area”, along with their position and scale.
Authoring Crowd Events ($Q2$). The second query focuses on defining the action sequence and interactions of agents, both with the environment and other agents. Specifically, this request takes as input $S_{in}$ and a simplified list of the previously generated environment. The output is a list of multiple crowd events. Each event contains a sequence of action-location pairs and the id of the agent/s that participate in the current event. Additionally, a trigger field defines the start condition of each event, e.g., E8 starts once E2 ends.
Finally, we combine the responses from both queries to construct a high-level crowd scenario $S$, based on a textual description $S_{in}$. For each agent $\alpha_{i}$} we record the sequence of events in which it appears and define a record $Rec_{\alpha_{i}}$. This record offers a comprehensive view of the agent’s behavior and interactions throughout its entire lifespan in the simulated environment.